OpenStudy (anonymous):

Two cards are drawn, without being replaced from a deck of 52 cards. Find the probability of selecting a king and a queen.

Parth (parthkohli):

\( \color{Black}{\Rightarrow P(\mathbb{KING}) \times P(\mathbb{QUEEN}) }\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {4 \over 52} \times {4 \over 52} }\) \( \color{Black}{\Rightarrow \Large {1 \over 13} \times {1 \over 13} }\) \( \color{Black}{\Rightarrow \Large {1 \over 13^2} }\)

OpenStudy (anonymous):

I don't know what you did

Parth (parthkohli):

What is the probability of selecting a king?

OpenStudy (anonymous):

of selecting a king and a queen

Parth (parthkohli):

No, first tell me what is the probability of selecting a king if you pick on card?

Parth (parthkohli):

one*

OpenStudy (anonymous):

idk...

Parth (parthkohli):

There are 4 kings....there are total 52. P(Event) = Possible number of outcomes/Total number of outcomes.

OpenStudy (unklerhaukus):

there are thirteen cards from each suit in the deck

OpenStudy (mertsj):

4/52 times 4/51

Parth (parthkohli):

Oh I see...seems like I missed the part "not being replaced"

OpenStudy (phi):

4/52 * 4/51 is prob of K followed by Q now what about Q followed by K?